Solve each of the following systems by substitution.
5x – 2y = -5
Y – 5x = 3

Answer by checkley71(8403)

Y=5X+3 OR SUB IN EQUATION #1 WE GET 5X-2(5X+3)=-5 OR 5X-10X-6=-5 OR -5X=1 OR
X=-1/5 THE 5(-1/5)-2Y=-5 OR -1-2Y=-5 OR -2Y=-4 OR Y=2